A circular region of radius R = 0.003 m has a uniform displacement current that flow perpendicular to the the plane with I = 0.005 A. What is the magnitude of the magnetic field due to the displacement current at a radial distance of 0.002 m and 0.005 m.
1
Expert's answer
2017-03-21T10:10:15-0400
Magnetic field of a circle can be calculated as B = mm0IR^2/2a^2. B1 = 6.04*10^-16 T and B2 = 1.43*10^-16 T.